OpenStudy (anonymous):

What is (6log_5 sqrt(4x-1) - log_5(9/x)+log_5 9 equal to ?

OpenStudy (anonymous):

\[6\log_{5}\sqrt{4x-1}-\log_5(\frac{9}{x})+\log_5 9\] ?

OpenStudy (anonymous):

yeep

OpenStudy (anonymous):

i don't have a graphing calc and the online one can't solve that =="

OpenStudy (anonymous):

if so take the coefficient from in front of the first term and apply it to the exponent of the log, this is a rule of logs: \[6\log_{5}(\sqrt{4x-1}) =\log_{5}(\sqrt{4x-1})^6\] \[=\log_{5}(4x-1)^3\]

OpenStudy (ksaimouli):

use web2.0 calculator

OpenStudy (anonymous):

with out a value for x you won't be able to solve it...i'd suggest simplify it as much as you can using log rules. ie. all the logs of similar base being added/subtracted can be combined